OpenStudy (anonymous):

In a survey conducted by a union, members were asked to rate the importance of the following issues: (1) job security, (2) increased fringe benefits, and (3) improved working conditions. Five different responses were allowed for each issue. Among completed surveys, how many different responses to this survey were possible?

OpenStudy (anonymous):

I know this needs to be done through Generalized multiplication principle but can't figure it out.

OpenStudy (anonymous):

(number of possible responses for issue 1) * (# for issue 2) * (issue 3) = 5 * 5 * 5 = 125

OpenStudy (anonymous):

oh wow..i was doing 5*3...

OpenStudy (anonymous):

Thank you so very much